OpenStudy (anonymous):

Solve. 3(x – 2) + 5x = 18 A. all real numbers B. 1 C. 3 D. 2

OpenStudy (solomonzelman):

\(\large\color{blue}{ \rm 3(x – 2) + 5x = 18 }\) \(\large\color{blue}{ \rm 3x – 6 + 5x = 18 }\) \(\large\color{blue}{ \rm 3x + 5x = 18+6 }\) \(\large\color{blue}{ \rm 8x = 24 }\) \(\large\color{blue}{ \rm x = ~~?? }\) you tell me .

OpenStudy (anonymous):

3?

OpenStudy (solomonzelman):

Yeah

OpenStudy (anonymous):

math magician this guy is.

OpenStudy (solomonzelman):

thanks kracker :)

OpenStudy (anonymous):

y u so racist?